OpenStudy (anonymous):

Type the ordered pair that is the solution to these equations. 3x - 2y = 8 2x + 5y = -1

OpenStudy (anonymous):

multiply first equation with 2 and second equation with 3 6x-4y=16 6x+15y=-3 subtract (new)first one from second one , this will eliminate x, -19y=19 y=-1 substitute y to get x in 3x - 2y = 8 3x+2=8 3x=6 x=2 answer is (2,-1) http://www.wolframalpha.com/input/?i=3x+-+2y+%3D+8%2C+2x+%2B+5y+%3D+-1